OpenStudy (anonymous):

how to find derivatives using the limit definition ?for f(x)=(1-3x)^1/2

OpenStudy (anonymous):

\[f(x)=(1-3x)^{1/2}\]

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \frac{ f(x+h)-f(x) }{ h }\] then i did this \[\lim_{h \rightarrow 0} \frac{ f(1-3(x+h)^{1/2}-(1-3x)^{1/2} }{ h }\] after that \[\lim_{h \rightarrow 0} \frac{ f(1-3x+3h)^{1/2}-(1-3x)^{1/2} }{ h }\] what is the next step to remove the "h" from bottom ?1

OpenStudy (anonymous):

rationalize numerator

OpenStudy (anonymous):

how can i do it?! @beccaboo022a

OpenStudy (anonymous):

\(\large \lim \limits_{h \rightarrow 0}~ \frac{ (1-3(x+h)^{1/2}-(1-3x)^{1/2} }{ h }\) \(\large \lim \limits_{h \rightarrow 0}~ \frac{ \sqrt{(1-3(x+h)}-\sqrt{(1-3x)} }{ h }\)

OpenStudy (anonymous):

multiply numerator and denominator wid conjugate of numerator

OpenStudy (anonymous):

\(\large \lim \limits_{h \rightarrow 0}~ \frac{ (1-3(x+h)^{1/2}-(1-3x)^{1/2} }{ h }\) \(\large \lim \limits_{h \rightarrow 0}~ \frac{ \sqrt{(1-3(x+h)}-\sqrt{(1-3x)} }{ h }\) \(\large \lim \limits_{h \rightarrow 0}~ \frac{ \sqrt{(1-3(x+h)}-\sqrt{(1-3x)} }{ h } \times \frac{\sqrt{(1-3(x+h)}+\sqrt{(1-3x)}}{\sqrt{(1-3(x+h)}+\sqrt{(1-3x)}}\)

OpenStudy (anonymous):

see something canncelling out ? :)

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \frac{1-3(x+h)-(1-3x) }{ h \sqrt{1-3(x+h)}-\sqrt{1-3x} }\] then \[\lim_{h \rightarrow 0} \frac{1-3x+3h-1+3x }{ h \sqrt{1-3(x+h)}-\sqrt{1-3x} }\] after that \[\lim_{h \rightarrow 0} \frac{3h }{ h \sqrt{1-3(x+h)}-\sqrt{1-3x} }\] lastly \[\lim_{h \rightarrow 0} \frac{3 }{ \sqrt{1-3(x)}-\sqrt{1-3x} }\] and finally \[\frac{3 }{ \sqrt{1-3x}-\sqrt{1-3x} }\] is it true @beccaboo022a

OpenStudy (anonymous):

nope. check ur signs again, check, conjugate is + sign...

OpenStudy (anonymous):

\(\large \lim \limits_{h \rightarrow 0}~ \frac{ \sqrt{(1-3(x+h)}-\sqrt{(1-3x)} }{ h } \times \frac{\sqrt{(1-3(x+h)}\color{red}{+}\sqrt{(1-3x)}}{\sqrt{(1-3(x+h)}\color{red}{+}\sqrt{(1-3x)}}\)

OpenStudy (anonymous):

\(\large \lim \limits_{h \rightarrow 0}~ \frac{ \sqrt{(1-3(x+h)}-\sqrt{(1-3x)} }{ h } \times \frac{\sqrt{(1-3(x+h)}\color{red}{+}\sqrt{(1-3x)}}{\sqrt{(1-3(x+h)}\color{red}{+}\sqrt{(1-3x)}}\) \(\large \lim \limits_{h \rightarrow 0}~ \frac{ (\sqrt{(1-3(x+h)})^2-(\sqrt{(1-3x)})^2 }{ h(\sqrt{(1-3(x+h)}\color{red}{+}\sqrt{(1-3x)}) } \) \(\large \lim \limits_{h \rightarrow 0}~ \frac{ 1-3(x+h)-(1-3x) }{ h(\sqrt{(1-3(x+h)}\color{red}{+}\sqrt{(1-3x)}) } \) \(\large \lim \limits_{h \rightarrow 0}~ \frac{ 1-3x-3h-1+3x }{ h(\sqrt{(1-3(x+h)}\color{red}{+}\sqrt{(1-3x)}) } \) \(\large \lim \limits_{h \rightarrow 0}~ \frac{ -3h }{ h(\sqrt{(1-3(x+h)}\color{red}{+}\sqrt{(1-3x)}) } \)

OpenStudy (anonymous):

keep going..

OpenStudy (anonymous):

ok thank you very much @beccaboo022a for showing steps .

OpenStudy (anonymous):

np :) let me knw the final answer

OpenStudy (anonymous):

OK \[\lim_{h \rightarrow 0} \frac{-3 }{ \sqrt{1-3(x+h)+\sqrt{1-3x}} }\] then \[\lim_{h \rightarrow 0} \frac{-3 }{ \sqrt{1-3x+3(0)+\sqrt{1-3x}} }\] == \[\frac{-3 }{ \sqrt{1-3x+\sqrt{1-3x}} }\] finally \[\frac{-3 }{ \sqrt{1-3x+\sqrt{1-3x}} }\] \[\frac{-3 }{ 2\sqrt{1-3x} } \] @beccaboo022a this it the final answer

OpenStudy (anonymous):

thats right ! good work :) just one notational mistake u were doing in all the limit problems

OpenStudy (anonymous):

\(\large \color{red}{\lim \limits_{h \rightarrow 0} }\frac{-3 }{ \sqrt{1-3x+3(0)+\sqrt{1-3x}} } \)

OpenStudy (anonymous):

you should not show that limit symbol after taking the limit / at the time of evaluating the limit

OpenStudy (anonymous):

thank you @beccaboo022a ,now i know my mistake i will never do it again .i am so happy to understand how to solve this problem .

OpenStudy (anonymous):

glad to hear that :))